No Streak, Inc. manufactures windshield wipers for its single client, Go Fast Motors. No Streak sells each wiper to Go Fast for $25. Direct material costs per unit equal $4 and direct labor cost per unit equals $1. Also, No Streak incurs $5 of variable overhead costs for every unit that it manufactures and sells. Finally, the fixed costs associated with No Streak’s manufacturing plant equal $20,000 and the fixed costs related to No Streak’s marketing and distribution equal $30,000.

1. [Basic breakeven analysis.] Calculate No Streak’s breakeven point in both output volume and sales dollars.

2. [Profit goals and taxes.] What output volume would No Streak need to achieve to generate a $15,000 pre-tax profit? What output volume would No Streak need to achieve to generate an after-tax profit of $15,000? [No Streak’s tax rate is 40%].

3. [Changing cost structures.] No Streak realizes that its biggest competitor, Spot Free, manufactures its windshield wipers in an overseas production plant. Spot Free is able to produce wipers at a significantly lower cost due to the lower labor cost in its overseas plant. In order to compete more effectively on costs, No Streak is considering changing its cost structure by automating a greater portion of its manufacturing process. Specifically, No Streak’s new cost structure would increase total fixed costs to $100,000 and reduce variable cost per unit to $5. At what level of output volume would No Streak be indifferent between the current cost structure and the proposed new cost structure? What profit is generated at this output volume? 0.5 pts

4. [Changing revenue and cost structures.] Assume the same information as in #3. However, No Streak believes that in addition to the new cost structure, it will also need to lower its price per unit to $23 to remain competitive. At what level of output volume would No Streak be indifferent between the current cost and revenue structure and this new alternative cost and revenue structure? What profit is generated at this output volume?

5. [Putting it all together.]

a. Which cost structure should No Streak select when comparing the status quo to the first alternative (i.e., in #3), or the status quo to the second alternative (i.e., in #4)? What factor plays a very large role in helping to answer this question?

b. Assume that No Streak’s production and marketing departments inform you that they expect to produce and sell 14,000 wipers, respectively. Given this information, do you recommend that No Streak should stay with the status quo OR implement the first alternative? Briefly explain your response.

c. After an unanticipated downturn in the economy, assume now that No Streak’s production and marketing departments inform you that they expect to produce and sell only 6,000 wipers, respectively. Does this reduction change your recommendation regarding the status quo versus the first alternative? Briefly explain your response.

d. Construct a Contribution Margin Income Statement that incorporates your recommended cost structure for No Streak at an output level of 6,000 units.

6. [Multiple Product CVP.] No Streak has decided to stay with its current cost and revenue structure. Instead, No Streak has elected to add Sheer Shine, a miraculous windshield cleaning solution, to its product mix offering. The variable cost to manufacture and sell each bottle of Sheer Shine is $7 and each bottle sells for $10. Additional fixed costs of $10,000 are required to manufacture Sheer Shine. Based on analyses of competitors, No Streak estimates that it would manufacture and sell two bottles of Sheer Shine for every three windshield wipers that it manufactures and sells. Assuming this product mix is accurate, how many windshield wipers and bottles of Sheer Shine must No Streak sell in order to breakeven this year? Also, if No Streak expects to sell 6,000 wipers and 4,000 Sheer Shine bottles, compute the Margin of Safety in units, sales dollars, and contribution margin.

7. [Sensitivity Analysis.] No streak defines CVP risk as the variation (or range) in net income that would result from specific changes in critical CVP inputs. Assume that No Streak expects to sell 6,000 Windshield Wipers and 4,000 bottles of Sheer Shine. Compute and rank in descending order of CVP risk each of the following three changes in CVP inputs associated with the multiple product CVP analysis in requirement 6: (A) The price for each bottle of Sheer Shine falls from $10 to $8. (B) The variable overhead cost for each windshield wiper increases from $5 to $10. (C) The fixed marketing and distribution costs increase from $30,000 to $45,000.
Answer #1

Selling price per unit = $25

Direct Meterial per unit = $4

Direct Labour per unit = $1

Variable overhead per unit = $5

Fixed Cost (Manufacturing) = $20,000

Fixed Cost (Marketing and Distribution) = $30,000

Contribution per unit = Selling price per unit - Variable cost per unit

Contribution per unit = $25 - ($4+$1+$5) = $15

PV Ratio = (Contribution per unit / Selling price per unit)*100 = ($15 / $25)*100 = 60%

Break Even Point = Fixed Cost / Contribution per Unit

Break Even Point = ($20,000+$30,000) / $15 = 3,333 Units

Sales to get $15,000 Profit (Pre Tax),

Sales in $ = (fixed cost + Profit) / PV Ratio = (50,000+15,000) / 60% = $108,333

Sales to get $15,000 Profit (Post Tax)

Tax Rate = 40% (Given in the Question)

Pre tax profit = 15000*100/(100-40) = $25000

Sales in $ = (fixed cost + Profit) / PV Ratio = (50,000+25,000) / 60% = $125,000

When,

Fixed Cost = $100,000

Variable Cost per unit = $5

Sales per unit = $25

Contribution per unit = Selling price per unit - Variable cost per unit

Contribution per unit = $25 - $5 = $20

PV Ratio = (Variable Cost per unit / Sales per unit)*100 = ($5 / $25)*100 = 80%

Break Even Point = Fixed Cost / PV Ratio

Break Even Point = ($100,000) / 80% = $125,000

When Selling price per unit = $23 and to achieve a total sales of $125,000

The no:of unit to be sold = $125,000/$23 = 5435 units